10x-(3-x)= 2(x+3)+ 9. The teacher says the answer is x=2, but i must show how to get that answer! please help!
Answer by stanbon(48569)

10x-(3-x)= 2(x+3)+ 9
Apply the distributive property to get rid of the parentheses.
10x-3+x=2x+6+9
Combine Like terms to get:
9x=18
Divide by 9 to get:
x=2
Cheers,
Stan H.